Complete the following chemical equations :
(a) KI + CI2 →      (b) KcIO3­ + I2
Justify the formation of the products in the above reactions.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
(a) 2KI + CI2 → 2KCI + I2
Since CI2 is more powerful oxidizing agent than I2, CI­2, is able to displace I- to form I2.
2I- → I2 + 2e-, E0 = + 0.54 V ….(i)
CI2 + 2e- → 2CI- E0 = 1.36V ….(ii)
On subtracting eq. (i) from eq. (ii), we get
CI_{2_{g}} + 2I_{(aq)}^{-} +\rightarrow 2CI_{(aq)}^{-} + I_{2_{(s)}} E^{0} = 82V
(b) 2KCIO3 + I2 → 2KIO3 + CI2
Here CIO_{3}^{-}is more powerful oxidizing agent than IO_{3}^{-}, so CI is displaced by I.
2IO_{3}^{-}+ 12H+ + 10e- → I2 + 6H2O, E0 = 1.195V ….(i)
2CIO_{3}^{-}+ 12H+ + 10e- → CI2 + 6H2O, E0 = 1.47V ….(ii)
On substracting eq. (i) from eq. (ii), we get
2CIO_{3}^{-}+ I2 2IO_{3}^{-}+ CI2 E0 = 0.275V